OpenStudy (anonymous):

The volume of a cube is found by using the formula L^3, where L is the side length. If the side length is 3x^2 what is the volume of the cube?

jhonyy9 (jhonyy9):

so then V=L^3 and you know that L=3x^2 then this mean that V = (3x^2)^3 = can you calcule it sure easy yes ?

jhonyy9 (jhonyy9):

V = 3^3 * (x^2)^3 = 27x^6 hope this will help you